\begin{thebibliography}{10}\itemsep=-1pt

\bibitem{bodla2017soft}
Navaneeth Bodla, Bharat Singh, Rama Chellappa, and Larry~S Davis.
\newblock {Soft-NMS--improving object detection with one line of code}.
\newblock In {\em ICCV}, 2017.

\bibitem{dai2017deformable}
Jifeng Dai, Haozhi Qi, Yuwen Xiong, Yi Li, Guodong Zhang, Han Hu, and Yichen
  Wei.
\newblock {Deformable Convolutional Networks}.
\newblock In {\em ICCV}, 2017.

\bibitem{fang2021read}
Shancheng Fang, Hongtao Xie, Yuxin Wang, Zhendong Mao, and Yongdong Zhang.
\newblock {Read Like Humans: Autonomous, Bidirectional and Iterative Language
  Modeling for Scene Text Recognition}.
\newblock In {\em CVPR}, 2021.

\bibitem{guo2022visual}
Meng-Hao Guo, Cheng-Ze Lu, Zheng-Ning Liu, Ming-Ming Cheng, and Shi-Min Hu.
\newblock {Visual Attention Network}.
\newblock {\em arXiv:2202.09741}, 2022.

\bibitem{Gupta16}
Ankush Gupta, Andrea Vedaldi, and Andrew Zisserman.
\newblock {Synthetic Data for Text Localisation in Natural Images}.
\newblock In {\em CVPR}, 2016.

\bibitem{karatzas2015icdar}
Dimosthenis Karatzas, Lluis Gomez-Bigorda, Anguelos Nicolaou, Suman Ghosh,
  Andrew Bagdanov, Masakazu Iwamura, Jiri Matas, Lukas Neumann,
  Vijay~Ramaseshan Chandrasekhar, Shijian Lu, et~al.
\newblock {ICDAR 2015 competition on Robust Reading}.
\newblock In {\em ICDAR}, 2015.

\bibitem{karatzas2013icdar}
Dimosthenis Karatzas, Faisal Shafait, Seiichi Uchida, Masakazu Iwamura,
  Lluis~Gomez i Bigorda, Sergi~Robles Mestre, Joan Mas, David~Fernandez Mota,
  Jon~Almazan Almazan, and Lluis~Pere De~Las~Heras.
\newblock {ICDAR 2013 Robust Reading Competition}.
\newblock In {\em ICDAR}, 2013.

\bibitem{krylov2021open}
Ilya Krylov, Sergei Nosov, and Vladislav Sovrasov.
\newblock {Open images v5 text annotation and yet another mask text spotter}.
\newblock In {\em ACML}, 2021.

\bibitem{liao2020mask}
Minghui Liao, Guan Pang, Jing Huang, Tal Hassner, and Xiang Bai.
\newblock {Mask TextSpotter v3: Segmentation Proposal Network for Robust Scene
  Text Spotting}.
\newblock In {\em ECCV}, 2020.

\bibitem{litman2020scatter}
Ron Litman, Oron Anschel, Shahar Tsiper, Roee Litman, Shai Mazor, and R
  Manmatha.
\newblock {SCATTER: Selective Context Attentional Scene Text Recognizer}.
\newblock In {\em CVPR}, 2020.

\bibitem{long2022towards}
Shangbang Long, Siyang Qin, Dmitry Panteleev, Alessandro Bissacco, Yasuhisa
  Fujii, and Michalis Raptis.
\newblock {Towards End-to-End Unified Scene Text Detection and Layout
  Analysis}.
\newblock In {\em CVPR}, 2022.

\bibitem{nayef2019icdar2019}
Nibal Nayef, Yash Patel, Michal Busta, Pinaki~Nath Chowdhury, Dimosthenis
  Karatzas, Wafa Khlif, Jiri Matas, Umapada Pal, Jean-Christophe Burie,
  Cheng-lin Liu, et~al.
\newblock {ICDAR2019 Robust Reading Challenge on Multi-lingual Scene Text
  Detection and Recognition -- RRC-MLT-2019}.
\newblock In {\em ICDAR}, 2019.

\bibitem{singh2021textocr}
Amanpreet Singh, Guan Pang, Mandy Toh, Jing Huang, Wojciech Galuba, and Tal
  Hassner.
\newblock {TextOCR: Towards large-scale end-to-end reasoning for
  arbitrary-shaped scene text}.
\newblock In {\em CVPR}, 2021.

\bibitem{veit2016coco}
Andreas Veit, Tomas Matera, Lukas Neumann, Jiri Matas, and Serge Belongie.
\newblock {COCO-Text: Dataset and Benchmark for Text Detection and Recognition
  in Natural Images}.
\newblock {\em arXiv:1601.07140}, 2016.

\bibitem{wang2019efficient}
Wenhai Wang, Enze Xie, Xiaoge Song, Yuhang Zang, Wenjia Wang, Tong Lu, Gang Yu,
  and Chunhua Shen.
\newblock {Efficient and accurate arbitrary-shaped text detection with pixel
  aggregation network}.
\newblock In {\em ICCV}, 2019.

\bibitem{xue2022language}
Chuhui Xue, Yu Hao, Shijian Lu, Philip Torr, and Song Bai.
\newblock {Language Matters: A Weakly Supervised Vision-Language Pre-training
  Approach for Scene Text Detection and Spotting}.
\newblock In {\em ECCV}, 2022.

\bibitem{Zhang_2022_CVPR}
Xiang Zhang, Yongwen Su, Subarna Tripathi, and Zhuowen Tu.
\newblock {Text Spotting Transformers}.
\newblock In {\em CVPR}, 2022.

\bibitem{zhu2022oovrec}
Zhangzi Zhu, Yu Hao, Wenqing Zhang, Chuhui Xue, and Song Bai.
\newblock {Runner-Up Solution to ECCV 2022 Challenge on Out of Vocabulary Scene
  Text Understanding: Cropped Word Recognition}.
\newblock {\em arXiv:2208.02747}, 2022.

\end{thebibliography}

\documentclass[10pt,twocolumn,letterpaper]{article}

\usepackage{cvpr}              

\usepackage{algorithm}
\usepackage{algorithmic}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{booktabs}
\usepackage{amsfonts}

\usepackage{multirow}
\usepackage{array}
\usepackage{color}
\usepackage{subcaption}
\usepackage{colortbl}
\usepackage{arydshln}
\graphicspath{{images/}}
\usepackage{bbding}
\usepackage{makecell}
\usepackage{hhline}
\usepackage{bm}

\usepackage[pagebackref,breaklinks,colorlinks]{hyperref}

\usepackage[capitalize]{cleveref}
\crefname{section}{Sec.}{Secs.}
\Crefname{section}{Section}{Sections}
\Crefname{table}{Table}{Tables}
\crefname{table}{Tab.}{Tabs.}

\def\cvprPaperID{*****} 
\def\confName{CVPR}
\def\confYear{2022}

\begin{document}

\title{1st Place Solution to ECCV 2022 Challenge on Out of Vocabulary Scene Text Understanding: End-to-End Recognition of Out of Vocabulary Words}

\author{Zhangzi Zhu, Chuhui Xue, Yu Hao, Wenqing Zhang, Song Bai\\
ByteDance Inc.
}
\maketitle

\begin{abstract}
Scene text recognition has attracted increasing interest in recent years due to its wide range of applications in multilingual translation, autonomous driving, etc. In this report, we describe our solution to the Out of Vocabulary Scene Text Understanding (OOV-ST) Challenge, which aims to extract out-of-vocabulary (OOV) words from natural scene images. Our oCLIP-based model achieves 28.59\% in h-mean which ranks 1st in end-to-end OOV word recognition track of OOV Challenge in ECCV2022 TiE Workshop. 
\end{abstract}

\section{Introduction}
Recently, many scene text recognition techniques have been proposed which learns the language knowledge for better recognition the in-vocabulary (IV) words (\textit{i.e.}, the words have been appeared in the training set). However, the language information of out-of-vocabulary (OOV) words are usually difficult to learn if they have never been seen during training, which makes model difficult to recognize the OOV words accurately.

ECCV 2022 Challenge on Out of Vocabulary Scene Text Understanding \footnote{\scriptsize\url{https://rrc.cvc.uab.es/?ch=19}}, held together with ECCV 2022 workshop on Text in Everything (TiE) \footnote{\scriptsize\url{https://sites.google.com/view/tie-eccv2022/challenge}},  aims to evaluate the model performances on recognizing OOV words. In this challenge, the training, validation and test sets are composed of several commonly used datasets, including ICDAR13 \cite{karatzas2013icdar}, ICDAR15 \cite{karatzas2015icdar}, MLT19 \cite{nayef2019icdar2019}, COCO-Text \cite{veit2016coco}, TextOCR \cite{singh2021textocr}, HierText \cite{long2022towards}, and OpenImagesText \cite{krylov2021open}. Two evaluation metrics are provided that focuses on: (1) OOV words only which aims to evaluate the model performances on recognizing the OOV words; and (2) both IV and OOV words by averaging the IV and OOV scores which aims to consider both IV and OOV words in evaluation.

In this report, we present our solution to the end-to-end OOV word recognition task. We first pre-train different commonly-used network backbones by using oCLIP \cite{xue2022language}. We then fine-tune PAN \cite{wang2019efficient}, Mask TextSpotter-v3 (MTS-v3) \cite{liao2020mask} and TESTR \cite{Zhang_2022_CVPR} on the composed datasets for word detection. Finally, we recognize the detected words by using a SCATTER-based \cite{litman2020scatter} recognizer. Our method achieves 28.59\% h-mean, which ranks 1st on end-to-end recognition.

\section{Methods and Experimental Results}

\subsection{Text Detection}
In our solution, we first pre-train different backbones including VAN-large \cite{guo2022visual} and Deformable ResNet-101 \cite{dai2017deformable} by using oCLIP \cite{xue2022language}. Next, we fine-tune PAN \cite{wang2019efficient}, MTS-v3 \cite{liao2020mask} and TESTR \cite{Zhang_2022_CVPR} on the composed datasets by using the pre-trained models. Finally, we combine the detection results from different models together and recognize the words following \cite{zhu2022oovrec}. All models are evaluated on the validation set of the composed dataset.

\subsubsection{Model Pre-train}
We first pre-train VAN-large \cite{guo2022visual} and Deformable ResNet-101 \cite{dai2017deformable} by using oCLIP \cite{xue2022language} on SynthText \cite{Gupta16} dataset as well as the provided composed dataset. Next, we fine-tune PAN \cite{wang2019efficient} (with VAN-large as backbone), MTS-v3 \cite{liao2020mask} (with Deformable ResNet-101 as backbone), and TESTR \cite{Zhang_2022_CVPR} (with Deformable ResNet-101 as backbone) on the composed dataset by using the pre-trained backbone weights. By pre-training using oCLIP, the performances of different models have been improved by 1\% to 3\% in Fscore as shown in Table \ref{tab:oclip}. 

\begin{table*}[!h]
\centering
\begin{tabular}{lcclccl}
\toprule
\multirow{2}{*}{\textbf{Method}} & \multicolumn{3}{c}{\textbf{OOV}} & \multicolumn{3}{c}{\textbf{All}} \\
\cmidrule(lr){2-4}  \cmidrule(l){5-7}
& \textbf{Precision} & \textbf{Recall} & \textbf{Fscore} & \textbf{Precision} & \textbf{Recall} & \textbf{Fscore} \\
\midrule
PAN                     & 65.36  & 68.71  & 67.00 & 83.36  & 56.18  & 67.21 \\
PAN+oCLIP               & 64.03  & 73.11  & \textbf{68.27 (+1.27)} & 83.37  & 61.64  & \textbf{70.88 (+3.67)} \\
MTS-v3                     & 77.13  & 48.31  & 59.41 & 87.61  & 42.16  & 56.93 \\
MTS-v3+oCLIP               & 77.55  & 48.83  & \textbf{59.93 (+0.52)} & 87.73  & 43.09  & \textbf{57.87 (+0.94)} \\
TESTR                   & 69.55  & 55.12  & 61.50 & 84.75  & 52.34  & 64.71 \\
TESTR+oCLIP             & 71.47  & 56.22  & \textbf{62.93 (+1.43)} & 85.93  & 55.83  & \textbf{65.73 (+1.02)} \\
\bottomrule
\end{tabular}
\caption{Text detection results by adopting oCLIP\cite{xue2022language} for backbone pre-training.}
\label{tab:oclip}
\end{table*}

\subsubsection{Model Ensemble}
Next, we collect the detection results from different models with different scales of images (\textit{i.e.}, 512, 960, 1280, 1600) as input which are hence combined together. We further apply soft-nms \cite{bodla2017soft} on the combined results and filter the detected boxes by a threshold of 0.92. Table \ref{tab:ensemble} shows the model ensemble results.

\begin{table}[!t]
\centering
\begin{tabular}{cccc}
\toprule
\multirow{2}{*}{\textbf{Method}} & \multicolumn{3}{c}{\textbf{OOV}}                       \\  
 \cmidrule(l){2-4}                                & \textbf{Precision} & \textbf{Recall} & \textbf{Fscore} \\
\midrule
PAN                              & 64.03  & 73.11  & 68.27                 \\
MTS-v3                           & 77.55  & 48.83  & 59.93                 \\
TESTR                            & 71.47  & 56.22  & 62.93                 \\
\midrule
Ensemble                     & 69.85              & 76.20           & 72.89           \\
\bottomrule
\end{tabular}
\caption{Text detection results by model ensemble.}
\label{tab:ensemble}
\end{table}

\begin{table}[!h]
\centering
\begin{tabular}{cccc}
\toprule
\multirow{2}{*}{\textbf{Set}} & \multicolumn{3}{c}{\textbf{OOV}}                       \\
\cmidrule(l){2-4}                                 & \textbf{Precision} & \textbf{Recall} & \textbf{Fscore} \\
\midrule
Validation         & 41.08              & 41.73           & 41.40          \\          
Test                      & 20.28              & 48.42           & 28.59           \\

\bottomrule
\end{tabular}
\caption{End-to-end recognition results on validation and test set, respectively.}
\label{tab:e2e}
\end{table}

\subsection{End-to-End Word Recognition}
We pass the detected texts to our recognition model \cite{zhu2022oovrec} and filter out the words that are recognized to be `ignore' texts to obtain the text recognition results. Table \ref{tab:e2e} shows the end-to-end recognition results from our models. 

\section{Conclusion}
This report presents our solutions to the end-to-end OOV word recognition task of ECCV 2022 Challenge on OOV-ST. We adopt oCLIP for model pre-train and model ensemble for better detection of texts in scenes. The presented solution ranks first in the end-to-end recognition of out of vocabulary words in the ECCV 2022 Challenges on OOV-ST.

{\small
\bibliographystyle{ieee_fullname}
\bibliography{egbib}
}

\end{document}